OpenStudy (aravindg):

if[1-tanx ][1-tanx]=2 then general value of tan x =???

OpenStudy (anonymous):

this looks more doable

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

first lets solve \[(1-z)(1-z)=2\] \[1-2z+z^2=2\] \[z^2-2z-1=0\] \[z=1\pm\sqrt{2}\]so \[\tan(x)=1\pm\sqrt{2}\]

OpenStudy (aravindg):

srry a mistake in question

OpenStudy (aravindg):

if[1-tan \[\theta\]][1-tanx]=2 then general value of [x+\[\theta\]]=???

OpenStudy (aravindg):

if[1-tan θ][1-tanx]=2 then general value of [x+θ]=???